Light of wavelength 1.7x10^-11m undergoes Compton scattering. What
is the kinetic energy, momentum magnitude, speed, and de Broglie wave-
length of the electron which is emitted when the light scatters through an
angle of 90degrees?
Answer #1

Compton scattering : \lambda_f = \lambda_i + \frac{h}{m_o c} (1 - \cos \theta)

\Rightarrow \lambda_f = 1.7* 10^{-11} + \frac{6.626 * 10^{-34}}{9.1 * 10^{-31} * 3* 10^8} (1 - \cos 90^o) =1.943 * 10^{-11} m

Kinetic Energy, K = \frac{h^2}{2 \lambda^2 m} = \frac{(6.626*10^{*-34})^2}{2 * (1.943 * 10^{-11})^2 * 9.1 * 10^{-31} } = 6.39 * 10^{-16} J

Momentum, p = \frac{h}{\lambda} = \frac{6.626 * 10^{-34}}{1.943 * 10^{-11}} = 3.41 * 10^{-23} kgm/s

Speed, v = \frac{p}{m} = \frac{ 3.41 * 10^{-23}}{9.1 * 10^{-31}} = 3.75 * 10^7 m/s

de Broglie Wavelength, \lambda_f = 1.943 * 10^{-11} m